OpenStudy (anonymous):

In the series (6,11,16,21), what is the 20th term?

OpenStudy (anonymous):

as you can see the difference between each term is 5 , so this is an arithmetic series

OpenStudy (anonymous):

using the formula \[a_n=a_1+(n-1)d\] where a1=6,n=20, and d =5 you get \[a_20=6+(20-1)5\]

OpenStudy (anonymous):

\[a_{20}=6+(19)(5)\]